OpenStudy (jonnywktn):

can someone help me with this im really stuck determine the fourier series for the function f (x) = {-9 when -pie < x < 0 {+9 when 0 < x < pie periodic over 2 pie

hartnn (hartnn):

Hi @jonnywktn, At which step are you stuck?

hartnn (hartnn):

can you show your steps, whatever you have tried till now?

OpenStudy (jonnywktn):

im not sure how to do it at all struggling to understand it was looking for help from start to finish if possible

OpenStudy (jonnywktn):

is this something you could help with please?

hartnn (hartnn):

yes, do you know the formulas? if not, please check this: http://www.sosmath.com/fourier/fourier1/fourier1.html especially from the part that says "This theorem helps associate a Fourier series to any $2\pi$-periodic function."

OpenStudy (jonnywktn):

all I have is what ive posted I was just looking for someone to complete the question so I could work through it as im struggeling to understand or follow notes online

hartnn (hartnn):

can you solve integrals of the type: \(\Large \int \limits_{-\pi }^{\pi } f(x) \cos nx \: dx\) ?

hartnn (hartnn):

I can provide you the entire solution, but you might not learn anything from it, you might just copy it...

OpenStudy (jonnywktn):

the next part it to change the limits to 8 and 0 so I need to work through the first part and then do the other so im not just going to copy it I would appreciate the entire solution and answer for the one posted so I can learn from that to enable me to complete the rest thanks

OpenStudy (jonnywktn):

are you able to do this please thank you

hartnn (hartnn):

The fourier series of f(x) is represented by \(\Large f(x) = a_0 + \sum \limits_{n=1}^\infty a_n \cos nx + b_n \sin nx\) The constants, \(a_0, a_n, b_n\) for the periodic function with period = \(2\pi\) can be calculated as : \(a_0 = \dfrac{1}{2\pi}\int \limits_{-\pi}^{\pi}f(x)dx\) \(a_n = \dfrac{1}{\pi}\int \limits_{-\pi}^{\pi}f(x) \cos nx dx\) \(b_n = \dfrac{1}{\pi}\int \limits_{-\pi}^{\pi}f(x) \sin nxdx\) here, f(x) = -9 \((-\pi < x < 0)\) +9 \((0 < x < \pi)\) \(~\\ \Large a_0 = \dfrac{1}{2\pi}\int \limits_{-\pi}^{\pi}f(x)dx \\ \Large = \dfrac{1}{2\pi}[\int \limits_{-\pi}^{0}f(x)dx+\int \limits_{0}^{\pi}f(x)dx ]\\ \Large = \dfrac{1}{2\pi}[\int \limits_{-\pi}^{0}-9dx+\int \limits_{0}^{\pi}9dx ] \\ \Large = \dfrac{1}{2\pi} [[-9x]^0_{{-\pi }}+[9x]_0^\pi ] \\ \Large = \dfrac{1}{2\pi}[0-9\pi +9\pi +0] \\ \Large \bbox[yellow ]{a_0=0} \)

hartnn (hartnn):

\(\Large a_n = \dfrac{1}{\pi}\int \limits_{-\pi}^{\pi}f(x) \cos nx dx \\ ~\\ \Large = \dfrac{1}{\pi}[\int \limits_{-\pi}^{0}f(x)\cos nx dx+\int \limits_{0}^{\pi}f(x)\cos nxdx ]\\ \Large = \dfrac{1}{\pi}[\int \limits_{-\pi}^{0}-9\cos nx dx+\int \limits_{0}^{\pi}9 \cos nx dx ] \\ \Large = \dfrac{1}{\pi} [[\dfrac{-9}{n}\sin nx]^0_{{-\pi }}+[\dfrac{9}{n}\sin nx]_0^\pi ] \\ \Large = \dfrac{1}{n\pi}[0-9\sin n\pi +9\sin n\pi +0] \\ \Large Since, \sin n\pi = 0\\ \Large \bbox[yellow ]{a_n=0} \) Since, f(x) is an odd function, we got \(a_0 =0, a_n =0\). [Sometimes, we can use this directly, for odd functions, a0, an are 0, for even functions, bn =0]

hartnn (hartnn):

\(\Large b_n = \dfrac{1}{\pi}\int \limits_{-\pi}^{\pi}f(x) \sin nx dx \\ ~\\ \Large = \dfrac{1}{\pi}[\int \limits_{-\pi}^{0}f(x)\sin nx dx+\int \limits_{0}^{\pi}f(x)\sin nxdx ]\\ \Large = \dfrac{1}{\pi}[\int \limits_{-\pi}^{0}-9\sin nx dx+\int \limits_{0}^{\pi}9 \sin nx dx ] \\ \Large = \dfrac{1}{\pi} [[\dfrac{9}{n}\cos nx]^0_{{-\pi }}+[\dfrac{-9}{n}\cos nx]_0^\pi ] \\ \Large = \dfrac{9}{n\pi}[1-\cos (-n\pi) -\cos n\pi +1] \\ \Large = \dfrac{9}{n\pi}[2-2\cos (n\pi)]\\ \Large Since, \cos n\pi = (-1)^n\\ \Large \bbox[yellow ]{b_n=\dfrac{18}{n\pi}[1-(-1)^n]\: } \)

hartnn (hartnn):

Hence, the fourier series of f(x) is \(\Large f(x) = a_0 + \sum \limits_{n=1}^\infty a_n \cos nx + b_n \sin nx \\ \Huge f(x) = \sum \limits_{n=1}^\infty \dfrac{18}{n\pi}(1-(-1)^n) \sin nx \) Thats your final answer :)

hartnn (hartnn):

you can express it as \(\large f(x) = \dfrac{36}{\pi} [\sin x + \dfrac{\sin 3x}{3}+\dfrac{\sin 5x}{5}+...]\)

hartnn (hartnn):

Let me know if you have any doubts @jonnywktn :)

OpenStudy (jonnywktn):

sorry for the late reply I think I follow that part but I have been asked to work out b when n = 1,2,3,4 & 5 would this be something you could help with note I managed to complete this part for the other question so thanks for your help on that bit its much appreciated

hartnn (hartnn):

sure, \(\Large b_n=\dfrac{18}{n\pi}[1-(-1)^n]\) n=1, (-1)^n = -1 , 1-(-1) = 1+1= 2 so b1 = 36/1\(\pi\) = 36/\(\pi\) got this?

OpenStudy (jonnywktn):

I think so could you show me n=5 just to confirm I understand

hartnn (hartnn):

going for n=2, (-1)^2 = 1 1-1 = 0 so, b2 = 0 doubts?? knowing how to get b1 and b2, can you please give a try to find others??

hartnn (hartnn):

sure, \(\Large b_n=\dfrac{18}{n\pi}[1-(-1)^n]\) n=5, (-1)^5 = -1 , 1-(-1) = 1+1= 2 so b5 = 36/5\(\pi\)

hartnn (hartnn):

in all, b2 = b4 = 0 b1 = 36/pi b3 = 36/3pi = 12/pi b5 = 36/5pi

OpenStudy (jonnywktn):

so would you divide out from that so the answer is 36/5pi or would you leave it as that?

hartnn (hartnn):

leave it as it is.